OpenStudy (morgengarner):

A van travels 300 miles on 15 gallons of gas. Find how many gallons the van needs to travel 360 miles. A. 72 gallons of gas B. 67 gallons of gas C. 18 gallons of gas D. 23 gallons of gas

OpenStudy (gnech):

This is math, but ill help you anyways

OpenStudy (gnech):

you have to do a proportion to figure this out

OpenStudy (gnech):

do you know how to do proportions?

OpenStudy (morgengarner):

oh sorry lol i put it in math then i for got what class i was doing lol and put it in enlish with out reading it again and kinda

OpenStudy (gnech):

so do you still need help? :)

OpenStudy (morgengarner):

yess

OpenStudy (gnech):

ok how do you set up a proportion?

OpenStudy (morgengarner):

okie ummm hold on

OpenStudy (morgengarner):

okie i don't no

OpenStudy (gnech):

you can travel 300 miles on 15 gallons of gas.. 300/15 how many gallons will the van need in order to travel 360 miles.. 360/x x=gallons van can travel at 360 miles

OpenStudy (gnech):

300 miles = 360 miles 15 gal x gal

OpenStudy (gnech):

pretend there's a line in between 300/15 and 360/x

OpenStudy (gnech):

in order to solve you have to cross multiply, then divide by 300 360*15=5400 5400/300=18 gallons of gas

OpenStudy (gnech):

here is another example of proportion if you don't understand: 36in is equal to 3ft. How many ft is equal to 60 in? 36 in = 60 in 3ft x ft 36/3 = 60/x 60*3=180 180/36 = 5 ft in 60 in

OpenStudy (gnech):

***5ft is equal to 60in

OpenStudy (samanthagreer):

300 divided by 15 is 20 so that mean the car can go 20 miles on 1 gallon so then multiply each of them till you get one that equals 360 20 * 18 =360 so your answer is 18 gallons of gas

OpenStudy (kendre):

All I have to say is Divide 300 by 15 and that should leave you very close to the answer...